OpenStudy (anonymous):

WILL FAN AND MEDAL what is the value of |-9+x|-32 when x=-5?

OpenStudy (anonymous):

@misty1212 Hi can you please help me with this homework question?:)

OpenStudy (anonymous):

@pooja195 Hi can you please help me with this hw question?:)

pooja195 (pooja195):

|−9−5|−32 =|−14|−32 =14−32 =?

OpenStudy (anonymous):

-18? @pooja195

pooja195 (pooja195):

yep :)

OpenStudy (anonymous):

Thanks!:) really helpful! @ pooja195